A student recorded the waiting time for an appointment. Here is the sample data set (in hours):

1.5        0.6          0.9            1.8         1.9       0.5        1.1       4.4        0.5        0.8

(d) What is the percentage falling within 1 standard deviation? Does it satisfy the Empirical Rule?

HTML EditorKeyboard Shortcuts
Answer #1

Here

sample mean

sample standard deviation

mean - 1*standard deviation = 1.4 - 1.173 = 0.227

mean + 1*standard deviation = 1.4 +1.173 = 2.573

Number of observations falling within 1 standard deviation, i.e. in the interval(0.227, 2.573) is = 9

The percentage falling within 1 standard deviation

According to empirical rule, 68% values fall within 1 standard deviation.